A body start from rest and moves uniformly or accelerate uniformly at 20m/sec for 4sec .it then continues with this velocity for another 10sec before it is uniformly brought to rest in a further 5sec .draw the velocity time graph for the motion or the body and find its acceleration and find its retadiation and distance travelled during the uniform velocity and total distance travelled

1
Expert's answer
2022-01-19T10:12:11-0500


(a) acceleration

"a_1=\frac{\Delta v_1}{\Delta t_1}=\frac{20-0}{4}=\rm 5\: m/s^2"

(b) retardation

"a_3=\frac{\Delta v_3}{\Delta t_3}=\frac{0-20}{5}=\rm -4\: m/s^2"

(c) distance travelled during the uniform velocity

"d_2=v_2\Delta t_2=20*10=200\:\rm m"

(d) total distance travelled is equal to area below graph

"d=\frac{10+19}{2}*20=290\:\rm m"